Soybean meal is 16% protein, cornmeal is 8% protein
How many pounds if each should be mixed together in order to get 320 lb mixture that is 15% protein.

Answer by rfer(16322)

soybeans =x=280 lbs
cornmeal=320-x=40 lbs
.16x+.08(320-x)=3320(.15)
.16x+25.6-.08x=48
.08x=22.4
x=280